### v1

# Non-perturbative Origin of Electroweak Scale via Higgs-portal: Dyson-Schwinger in Conformally Invariant Scalar Sector

Abstract
We investigate conformally extended Standard Model with a hidden scalar $\phi$. It is shown that due to non-perturbative dynamics in the hidden sector, $\phi$ develops a vacuum expectation value (vev) in the form of a mass gap which triggers the electroweak symmetry breaking (EWSB) and dynamically generates the SM Higgs boson mass. For estimating the non-perturbatively generated mass scale, we solve the hierarchy of Dyson-Schwinger Equations in form of partial differential equations using the exact solution known via a novel technique developed by Bender, Milton and Savage. We employ Jacobi Elliptic function as exact background solution and show that the mass gap that arises in the hidden sector can be transmuted to the EW sector, expressed in terms of Higgs-portal mixed quartic coupling $\beta$ and self interaction quartic coupling $\lambda_{\phi}$ of $\phi$. We identify the suitable parameter space where the observed SM Higgs boson can be successfully generated . Finally, we discuss how this idea of non-perturbative EW scale generation can serve as a new starting point for better realistic model building in the context of resolving the hierarchy problem in the Standard Model.
I Introduction
The presence of any fundamental scalar field in quantum field theory (QFT) encounters the hierarchy problem because of the fact that its mass should be of the same order as that of the cut-off scale of the theory due to quantum corrections. As an example, if Planck scale is the highest scale in the theory, the Standard Model (SM) Higgs, if it is a fundamental scalar, should receive quantum radiative corrections to its mass of order of Planck scale, where quantum gravity is thought to be dominant  [^1] However, recently in higher-derivative non-local extensions of QFT, inspired by p-adic string field theories, it has been shown that this problem can be relaxed, and conformal invariance can be dynamically achieved without introducing any new particles in the physical mass spectrum, see Refs. [56, 54, 58, 44, 42, 50] with very concrete predictions and interesting LHC phenomenology [9, 88]..
Among several possibilities including supersymmetry and extra-dimensional theories one popular and quite elegant solution to this problem is to hypothesize scale invariance as a symmetry of the fundamental action, such that all scales we observe in Nature be generated dynamically beyond classical level. Coleman and Weinberg [17] proposed such a scenario and showed that the SM gauge symmetry breaking could be radiatively triggered via quantum corrections. However, from the observational point-of-view this mechanism fails within the Standard Model to generate the correct Higgs mass (the Electroweak (EW) Scale) as it predicts $m_{Z,W}>m_{H}$, where $m_{Z,W}$ are the SM Z and W gauge boson masses while $m_{H}$ is the SM Higgs boson mass [17, 21]. Inspite of this faliure this has remained the direction of BSM model building and several extensions of the SM has been explored extensively in the literature [1, 17, 86, 18, 20, 19], where this radiative EW symmetry breaking (EWSB) mechanism is successful in terms of observations and often makes very concrete and testable predictions. When non-minimal coupling to gravity is introduced, such scenarios can provide naturally flat inflaton potentials [77, 74, 84, 86, 72, 71, 7, 89] and stable particle dark matter candidates [62, 75, 72, 73, 76, 6]. The scenarios also lead to very strong first-order phase transitions and hence the possibility of high amplitude detectable gravitational wave (GW) signals in upcoming detectors [70, 82, 69, 5, 83, 11, 81, 60]. Consequently, scale invariant scenarios offer an interesting direction of model-building for solving the hierarchy problem in the Standard Model of particle physics [24, 2, 21, 62, 22, 3, 65, 86, 18, 72, 23, 73]. See Refs. [63, 4, 66, 67, 68, 12, 10] for other studies of conformal invariance and dimensional transmutation of energy scales along similar lines [85, 13, 52, 45, 59, 43, 55, 57].
Note that all the above mentioned analysis considered only the weak perturbation theory. In this paper, we develop a novel method to investigate such dynamical Higgs mass generation due to non-perturbative dynamics. We particularly focus on an scalar extension of the SM with a hidden sector dynamics triggering the EWSB. We develop a novel technique to solve series of Dyson-Schwinger Equations using exact Green’s function solution of the background Equation of Motion involving Jacobi Elliptical function, following the analytic approach of Dyson-Schwinger equations which is originally devised by Bender, Milton and Savage in Ref. [8]. Since the approach remains valid even in the strongly-coupled regime [39], this technique has been recently applied to QCD in Refs. [46, 47, 48, 49, 15] and to the SM Higgs sector in Ref. [37], as well as to other types of SM extensions over the past two decades, see Refs. [41, 15, 40, 38, 39, 37, 36, 35, 34, 33, 32, 31, 30, 29, 28, 26, 27, 25]. Recently, these have been employed to study non-perturbative hadronic contributions to the muon anomalous magnetic moment (g-2)μ [46], QCD in the non-perturbative regime [47, 48, 49], non-perturbative false vacuum decay [53], as well as to explore the mass gap and confinement in string-inspired infinite-derivative and Lee-Wick theories [44, 42, 51].
In the literature, the very simple idea of dynamical generation of EW scale from strong interactions has been studied extensively [64, 80, 87], a scenario very well-known as techni-color in old days. In its earliest version, the EW interactions of the matter content, in the form of fermions like techni-quarks Q were postulated such that their condensates breaks the EW symmetry and the EW scale originates from the dynamical scale of the technicolor physics. Later on this scenario was found disfavored in term of the flavor observables, the EW precision data and the measurements of the Higgs properties after its discovery. Later on, alternative strong dynamics was invoked to generate a composite or partially-composite Higgs boson, which more or less had an approach of postulating effective Lagrangians than fully realistically complete models. Phenomenologically viable dynamical models leading to interesting LHC phenomenology have been studied in [78, 79], although these models do not break EW symmetry nor provide a composite Higgs.
Assuming that quadratically divergent corrections to the Higgs mass squared have no physical meaning and hence can be ignored, possibly because the fundamental theory does not contain any mass term, one may promote a scale-invariant [^2] We will be using “scale-invariance” and “conformal invariance” inter-changeably in our paper, since it has been shown that they are classically equivalent for any four-dimensional field theory which respects unitarity and renormalizability[61, 14, 16]. symmetry principle at the classical level. In this context, dynamical generation of the EW scale via dimensional transmutation can be realised in models where an extra scalar $\phi$ develps a vacuum expectation value (VEV) through non-perturbative from $\lambda|\phi|^{4}$ interaction and then its interaction $\beta H^{2}|\phi|^{2}$ effectively generates a negative Higgs mass squared $m^{2}\sim-\beta\langle\phi\rangle^{2}$ with $\beta>0$ for the SM Higgs doublet $H$.
In order to study the non-perturbative dynamics for $\lambda_{\phi}\phi^{4}$ interaction, we utilize the exact solutions found in terms of the Jacobi elliptical functions following the analytic approach of Dyson-Schwinger equations originally devised by Bender, Milton and Savage in Ref. [8]. In this case, the Green’s functions of the theory are represented analytically, and therefore it is straightforward to understand the effect of the background on the interactions that remain valid even in the strongly-coupled regime [39].
The paper is organized as follows: we start in the next section by a short review of the strongly-coupled technique to generate mass gap via solving Dyson-Schwinger Equations in terms of Jacobi Elliptic function. In section III, we will discuss a simple extension of the SM involving Higgs-portal extra scalar singlet $\phi$ and generate EWSB dynamically. Section IV is devoted to conclusions and discussions.
II Mass Gap via Jacobi Elliptic Functions: short review
As a starting point, one considers the partition function. E.g., for a theory with action $S[\phi]$ one has
$Z[j]={\cal N}\int[d\phi]e^{iS[\phi]-i\int d^{4}xj(x)\phi(x)}$ (1)
with a scalar field $\phi(x)$. It is obvious that this functional integral does not change after a re-parametrization $\phi(x)\rightarrow\phi(x)+\alpha(x)$, with an arbitrary function $\alpha(x)$. Therefore,
$Z[j]\rightarrow Z[j]\langle e^{i\int d^{4}x\alpha(x)\left(\frac{\delta S}{\delta\phi(x)}-j(x)\phi(x)\right)}\rangle_{j},$ (2)
from which, by requiring invariance, we derive the quantum equation of motion:
$\left\langle\frac{\delta S}{\delta\phi(x)}\right\rangle_{j}=j(x).$ (3)
Repeating the derivation with respect to the current $j$ will give all the full set of Dyson-Schwinger equations for the correlation functions after setting $j=0$ at the end of computation. We note that from the lhs of eq. (3) one gets the average on the classical equations of motion of the theory that are the starting point for the procedure.
Bender-Milton-Savage method [8] takes the move from eq. (3), working always with higher-order n-point functions $G_{n}(x_{1},x_{2},\ldots,x_{n})$, without explicitly introducing vertex parts. This permits to preserve the differential structure of the Dyson-Schwinger equations, making this approach particularly useful when exact solutions are known. E.g., this will give for a $\phi^{4}$ theory [39]
$\partial^{2}G_{1}+m^{2}G_{1}+\lambda G_{1}^{3}+3\lambda G_{2}(0)+\lambda G_{3}(x,x,x)=0,$
$(\partial^{2}+m^{2})G_{2}(x-y)+3\lambda G_{2}(0)G_{2}(x-y)+3\lambda[G_{1}(x)]^{2}G_{2}(x-y)$
$+3\lambda G_{3}(x,x,y)G_{1}(x)+\lambda G_{4}(x,x,x,y))=\delta^{4}(x-y).$ (4)
Exact and non-trivial solutions for $G_{1}$ are now known: for $G_{3}(x,x,x)=0$, the the set of Dyson-Schwinger equations becomes treatable without any truncation.
Next we will move onto an SM extension where this BSM sector will be able to generate EWSB dynamically.
III SM Higgs mass from hidden sector mass gap
Suppose the system has calssically conformal symmetry. Now the only possibility to write terms in 4-D with 2 scalar fields and strictly re-normalizable (in t’Hooft-Veltman sense) is the following Lagrangian:
$\mathcal{L}=\frac{1}{2}(\partial\phi)^{2}+\frac{1}{2}(\partial h)^{2}-\frac{\lambda_{\phi}}{4}\phi^{4}-\frac{\lambda_{h}}{4}h^{4}+\beta\phi^{2}h^{2},$ (5)
where $\phi$ is a scalar field, $h$ is the SM Higgs field, $\beta$ is their interaction coupling, $\lambda_{\phi}$ and $\lambda_{h}$ are the self-interaction couplings. We assume $\lambda_{h}\ll 1$ and $\beta\ll\lambda_{\phi}$.
III.1 Dyson-Schwinger equations
Classical equations of motion are easily obtained to be
$\partial^{2}\phi$ $=$ $-\lambda_{\phi}\phi^{3}+\beta h^{2}\phi+j_{\phi},$
$\partial^{2}h$ $=$ $-\lambda_{h}h^{3}+\beta\phi^{2}h+j_{h}.$ (6)
Given the partition function
$Z[j_{\phi},j_{h}]=\int[d\phi][dh]\exp\left[-\int d^{4}xL-\int d^{4}x(j_{\phi}\phi+j_{h}h)\right],$ (7)
we can evaluate the averages as
$\partial^{2}G_{1}^{\phi}$ $=$ $-\lambda_{\phi}Z^{-1}[j_{\phi},j_{h}]\langle\phi^{3}\rangle+\beta Z^{-1}[j_{\phi},j_{h}]\langle h^{2}\phi\rangle+j_{\phi}$
$\partial^{2}G_{1}^{h}$ $=$ $-\lambda_{h}Z^{-1}[j_{\phi},j_{h}]\langle h^{3}\rangle+\beta Z^{-1}[j_{\phi},j_{h}]\langle\phi^{2}h\rangle+j_{h}.$ (8)
We notice that
$G_{1}^{\phi}(x)Z[j_{\phi},j_{h}]=\langle\phi(x)\rangle$
$G_{2}^{\phi\phi}(x,x)Z[j_{\phi},j_{h}]+[G_{1}^{\phi}(x)]^{2}Z[j_{\phi},j_{h}]=\langle\phi^{2}(x)\rangle$
$G_{3}^{\phi\phi h}(x,x,x)Z[j_{\phi},j_{h}]+G_{2}^{\phi\phi}(x,x)G_{1}^{h}(x)Z[j_{\phi},j_{h}]+2G_{2}^{\phi h}(x,x)G_{1}^{\phi}(x)Z[j_{\phi},j_{h}]+$
$[G_{1}^{\phi}(x)]^{2}G_{1}^{h}(x)Z[j_{\phi},j_{h}]=\langle\phi^{2}(x)h(x)\rangle$ (9)
Similarly, interchanging $\phi$ with $h$ will yield
$G_{1}^{h}(x)Z[j_{\phi},j_{h}]=\langle h(x)\rangle$
$G_{2}^{hh}(x,x)Z[j_{\phi},j_{h}]+[G_{1}^{h}(x)]^{2}Z[j_{\phi},j_{h}]=\langle h^{2}(x)\rangle$
$G_{3}^{hh\phi}(x,x,x)Z[j_{\phi},j_{h}]+G_{2}^{hh}(x,x)G_{1}^{\phi}(x)Z[j_{\phi},j_{h}]+2G_{2}^{h\phi}(x,x)G_{1}^{h}(x)Z[j_{\phi},j_{h}]+$
$[G_{1}^{h}(x)]^{2}G_{1}^{\phi}(x)Z[j_{\phi},j_{h}]=\langle h^{2}(x)\phi(x)\rangle.$ (10)
Therefore, one has the equations for the 1P-functions for $\phi$ field
$\partial^{2}G_{1}^{\phi}(x)+\lambda_{\phi}\left\{[G_{1}^{\phi}(x)]^{3}+3G_{2}^{\phi\phi}(x,x)G_{1}^{\phi}(x)+G_{3}^{\phi\phi\phi}(x,x,x)\right\}=$
$\beta\left\{G_{3}^{hh\phi}(x,x,x)+G_{2}^{hh}(x,x)G_{1}^{\phi}(x)+2G_{2}^{h\phi}(x,x)G_{1}^{h}(x)+[G_{1}^{h}(x)]^{2}G_{1}^{\phi}(x)\right\}+j_{\phi},$ (11)
and for the Higgs field
$\partial^{2}G_{1}^{h}(x)+\lambda_{h}\left\{[G_{1}^{h}(x)]^{3}+3G_{2}^{hh}(x,x)G_{1}^{h}(x)+G_{3}^{hhh}(x,x,x)\right\}=$
$\beta\left\{G_{3}^{\phi\phi h}(x,x,x)+G_{2}^{\phi\phi}(x,x)G_{1}^{h}(x)+2G_{2}^{\phi h}(x,x)G_{1}^{\phi}(x)+[G_{1}^{\phi}(x)]^{2}G_{1}^{h}(x)\right\}+j_{h}.$ (12)
Setting to zero the 3P-function at the same point and the currents, one has for the 1P-functions
$\partial^{2}H_{1}^{\phi}(x)+\lambda_{\phi}\left\{[H_{1}^{\phi}(x)]^{3}+3H_{2}^{\phi\phi}(0)H_{1}^{\phi}(x)\right\}=$
$\beta\left\{H_{2}^{hh}(0)H_{1}^{\phi}(x)+2H_{2}^{h\phi}(0)H_{1}^{h}(x)+[H_{1}^{h}(x)]^{2}H_{1}^{\phi}(x)\right\},$ (13)
and for the Higgs field
$\partial^{2}H_{1}^{h}(x)+\lambda_{h}\left\{[H_{1}^{h}(x)]^{3}+3H_{2}^{hh}(0)H_{1}^{h}(x)\right\}=$
$\beta\left\{H_{2}^{\phi\phi}(0)H_{1}^{h}(x)+2H_{2}^{\phi h}(0)H_{1}^{\phi}(x)+[H_{1}^{\phi}(x)]^{2}H_{1}^{h}(x)\right\}.$ (14)
We further set $H_{2}^{h\phi}(0)=H_{2}^{\phi h}(0)=0$ and get the full symmetrical set
$\partial^{2}H_{1}^{\phi}(x)+\lambda_{\phi}[H_{1}^{\phi}(x)]^{3}+3\lambda_{\phi}H_{2}^{\phi\phi}(0)H_{1}^{\phi}(x)$
$-\beta H_{2}^{hh}(0)H_{1}^{\phi}(x)-\beta[H_{1}^{h}(x)]^{2}H_{1}^{\phi}(x)=0,$ (15)
and for the Higgs field
$\partial^{2}H_{1}^{h}(x)+\lambda_{h}[H_{1}^{h}(x)]^{3}+3\lambda_{h}H_{2}^{hh}(0)H_{1}^{h}(x)$
$-\beta H_{2}^{\phi\phi}(0)H_{1}^{h}(x)-\beta[H_{1}^{\phi}(x)]^{2}H_{1}^{h}(x)=0.$ (16)
Now, we can make an approximation that $\beta[H_{1}^{\phi}(x)]^{2}H_{1}^{h}(x)$ is small with respect to the other terms in the equation and the solution $H_{1}^{h}(x)=v$ with a constant $v$, holds at the leading order (mean field approximation). This will yield for the $\phi$ field
$\partial^{2}H_{1}^{\phi}(x)+\lambda_{\phi}[H_{1}^{\phi}(x)]^{3}+\mu^{2}_{\phi}H_{1}^{\phi}(x)=0,$ (17)
where
$\mu_{\phi}^{2}=3\lambda_{\phi}H_{2}^{\phi\phi}(0)-\beta H_{2}^{hh}(0)-\beta v^{2}.$ (18)
This equation can be solved exactly. Note that this approximation holds even if the $\phi$ field is strongly coupled. For consistency reason, one should have in eq.(III.1)
$3\lambda_{h}H_{2}^{hh}(0)-\beta H_{2}^{\phi\phi}(0)<0,$ (19)
for the Higgs sector. This grants the correct vacuum expectation value for the theory for small $\beta$.
III.2 Gap equations
We can solve eq.(17) and obtain the corresponding 2P-function of the form:
$H_{1}^{\phi}(x)=\sqrt{\frac{2\mu^{4}}{\mu_{\phi}^{2}+\sqrt{\mu_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}}{\rm sn}\left(p\cdot x+\chi,\kappa\right),$ (20)
where $\mu$ and $\chi$ are arbitrary integration constants, $\kappa=\frac{-\mu_{\phi}^{2}+\sqrt{\mu_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}{-\mu_{\phi}^{2}-\sqrt{\mu_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}$, and the momentum $p$ is given by
$p^{2}=\mu_{\phi}^{2}+\frac{\lambda_{\phi}\mu^{4}}{\mu_{\phi}^{2}+\sqrt{\mu_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}.$ (21)
We need the 2P-functions that can be obtained from eq.(III.1) and (III.1). Setting the currents to zero at the end of computation, one has
$\partial^{2}H_{2}^{\phi\phi}(x,y)+3\lambda_{\phi}[H_{1}^{\phi}(x)]^{2}H_{2}^{\phi\phi}(x,y)+3\lambda_{\phi}H_{2}^{\phi\phi}(x,x)H_{2}^{\phi\phi}(x,y)-\beta[H_{1}^{h}(x)]^{2}H_{2}^{\phi\phi}(x,y)+$
$+\lambda_{\phi}\left\{3H_{3}^{\phi\phi\phi}(x,x,y)H_{1}^{\phi}(x)+H_{4}^{\phi\phi\phi\phi}(x,x,x,y)\right\}=$
$\beta\left\{H_{4}^{hh\phi\phi}(x,x,x,y)+H_{3}^{hh\phi}(x,x,y)H_{1}^{\phi}(x)+H_{2}^{hh}(x,x)H_{2}^{\phi\phi}(x,y)+2H_{3}^{h\phi\phi}(x,x,y)H_{1}^{h}(x)+\right.$
$\left.2H_{2}^{h\phi}(x,x)H_{2}^{h\phi}(x,y)+2H_{2}^{h\phi}(x,y)H_{1}^{\phi}(x)\right\}+\delta^{4}(x-y),$ (22)
and for the Higgs field
$\partial^{2}H_{2}^{hh}(x,y)+3\lambda_{h}[H_{1}^{h}(x)]^{2}H_{2}^{hh}(x,y)+3\lambda_{h}H_{2}^{hh}(x,x)H_{2}^{hh}(x,y)-\beta[H_{1}^{\phi}(x)]^{2}H_{2}^{hh}(x,y)$
$+\lambda_{\phi}\left\{3H_{3}^{hhh}(x,x,y)H_{1}^{h}(x)+H_{4}^{hhhh}(x,x,x,y)\right\}=$
$\beta\left\{H_{4}^{\phi\phi hh}(x,x,x,y)+H_{3}^{\phi hh}(x,x,y)H_{1}^{h}(x)+H_{2}^{\phi\phi}(x,x)H_{2}^{hh}(x,y)+2H_{3}^{\phi hh}(x,x,y)H_{1}^{\phi}(x)+\right.$
$\left.2H_{2}^{\phi h}(x,x)H_{2}^{\phi h}(x,y)+2H_{2}^{\phi h}(x,y)H_{1}^{h}(x)\right\}+\delta^{4}(x-y).$ (23)
In order to simplify these equations, we set the 1P-function for the Higgs field to be $H_{1}^{h}(x)=v$. This yields
$\partial^{2}H_{2}^{\phi\phi}(x,y)+3\lambda_{\phi}[H_{1}^{\phi}(x)]^{2}H_{2}^{\phi\phi}(x,y)+3\lambda_{\phi}H_{2}^{\phi\phi}(x,x)H_{2}^{\phi\phi}(x,y)-\beta v^{2}H_{2}^{\phi\phi}(x,y)+$
$+\lambda_{\phi}\left\{3H_{3}^{\phi\phi\phi}(x,x,y)H_{1}^{\phi}(x)+H_{4}^{\phi\phi\phi\phi}(x,x,x,y)\right\}=$
$\beta\left\{H_{4}^{hh\phi\phi}(x,x,x,y)+H_{3}^{hh\phi}(x,x,y)H_{1}^{\phi}(x)+H_{2}^{hh}(x,x)H_{2}^{\phi\phi}(x,y)+2vH_{3}^{h\phi\phi}(x,x,y)+\right.$
$\left.2H_{2}^{h\phi}(x,x)H_{2}^{h\phi}(x,y)+2H_{2}^{h\phi}(x,y)H_{1}^{\phi}(x)\right\}+\delta^{4}(x-y),$ (24)
and for the Higgs field
$\partial^{2}H_{2}^{hh}(x,y)+3\lambda_{h}v^{2}H_{2}^{hh}(x,y)+3\lambda_{h}H_{2}^{hh}(x,x)H_{2}^{hh}(x,y)-\beta[H_{1}^{\phi}(x)]^{2}H_{2}^{hh}(x,y)$
$+\lambda_{\phi}\left\{3vH_{3}^{hhh}(x,x,y)+H_{4}^{hhhh}(x,x,x,y)\right\}=$
$\beta\left\{H_{4}^{\phi\phi hh}(x,x,x,y)+vH_{3}^{\phi hh}(x,x,y)+H_{2}^{\phi\phi}(x,x)H_{2}^{hh}(x,y)+2H_{3}^{\phi hh}(x,x,y)H_{1}^{\phi}(x)+\right.$
$\left.2H_{2}^{\phi h}(x,x)H_{2}^{\phi h}(x,y)+2vH_{2}^{\phi h}(x,y)\right\}+\delta^{4}(x-y).$ (25)
These equations can be simplified further if we observe that higher-order nP-functions evaluated at the same space-time points can be chosen to be zero. In this way, one has
$\partial^{2}H_{2}^{\phi\phi}(x,y)+3\lambda_{\phi}[H_{1}^{\phi}(x)]^{2}H_{2}^{\phi\phi}(x,y)+3\lambda_{\phi}H_{2}^{\phi\phi}(0)H_{2}^{\phi\phi}(x,y)-\beta v^{2}H_{2}^{\phi\phi}(x,y)=$
$\beta\left\{H_{2}^{hh}(0)H_{2}^{\phi\phi}(x,y)+\right.$
$\left.2H_{2}^{h\phi}(0)H_{2}^{h\phi}(x,y)+2H_{2}^{h\phi}(x,y)H_{1}^{\phi}(x)\right\}+\delta^{4}(x-y),$ (26)
and for the Higgs field
$\partial^{2}H_{2}^{hh}(x,y)+3\lambda_{h}v^{2}H_{2}^{hh}(x,y)+3\lambda_{h}H_{2}^{hh}(0)H_{2}^{hh}(x,y)-\beta[H_{1}^{\phi}(x)]^{2}H_{2}^{hh}(x,y)=$
$\beta\left\{H_{2}^{\phi\phi}(0)H_{2}^{hh}(x,y)+2H_{2}^{\phi h}(0)H_{2}^{\phi h}(x,y)+2vH_{2}^{\phi h}(x,y)\right\}+\delta^{4}(x-y).$ (27)
Therefore, we can introduce the Green functions as
$\partial^{2}G_{2}^{\phi\phi}(x,y)+3\lambda_{\phi}[H_{1}^{\phi}(x)]^{2}G_{2}^{\phi\phi}(x,y)+m_{\phi}^{2}G_{2}^{\phi\phi}(x,y)=\delta^{4}(x-y)$
$\partial^{2}G_{2}^{hh}(x,y)+m_{h}^{2}G_{2}^{hh}(x,y)=\delta^{4}(x-y),$ (28)
where we have introduced the mass shift for the $\phi$ field and the mass of the Higgs field as
$m_{\phi}^{2}$ $=$ $3\lambda_{\phi}H_{2}^{\phi\phi}(0)-\beta v^{2}-\beta H_{2}^{hh}(0),$
$m_{h}^{2}$ $=$ $3\lambda_{h}v^{2}+3\lambda_{h}H_{2}^{hh}(0)-\beta H_{2}^{\phi\phi}(0).$ (29)
In the following, we will assume $\beta$ as a small positive parameter. These form a set of two gap equations. Indeed, we can write the following solutions to eqs.(III.2) and (III.2)
$H_{2}^{\phi\phi}(x,y)$ $=$ $G_{2}^{\phi\phi}(x,y)+\beta\int d^{4}zG_{2}^{\phi\phi}(x,z)\left[2H_{2}^{h\phi}(0)H_{2}^{h\phi}(z,y)+2H_{2}^{h\phi}(z,y)H_{1}^{\phi}(z)\right]$
$H_{2}^{hh}(x,y)$ $=$ $G_{2}^{hh}(x,y)+\beta\int d^{4}zG_{2}^{hh}(x,z)\left[2H_{2}^{\phi h}(0)H_{2}^{\phi h}(z,y)+2vH_{2}^{\phi h}(z,y)\right].$ (30)
We just note that these are perturbative equations as the cross-correlation functions $H^{h\phi}$ and $H^{\phi h}$ depend on $H^{\phi\phi}$ and $H^{hh}$. The propagators can be written in the form
$G_{2}^{\phi\phi}(p)=M_{\phi}{\hat{Z}}(m_{\phi},\lambda_{\phi})\frac{2\pi^{3}}{K^{3}(\kappa)}\sum_{n=0}^{\infty}(-1)^{n}\frac{e^{-(n+\frac{1}{2})\pi\frac{K^{\prime}(\kappa)}{K(\kappa)}}}{1-e^{-(2n+1)\frac{K^{\prime}(\kappa)}{K(\kappa)}\pi}}(2n+1)^{2}\frac{1}{p^{2}-m_{n}^{2}+i\epsilon},$ (31)
where
$M_{\phi}=\sqrt{m_{\phi}^{2}+\frac{\lambda_{\phi}\mu^{4}}{m_{\phi}^{2}+\sqrt{m_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}},$ (32)
${\hat{Z}}(m_{\phi},\lambda_{\phi})$ is a given constant, and $\mu$ is an integration constant. The spectrum is given by
$m_{n}=(2n+1)\frac{\pi}{2K(\kappa)}M_{\phi},$ (33)
with $K(\kappa)$ being the complete elliptic integral of the first kind.
For the Higgs field, we have
$G_{2}^{hh}(p)=\frac{1}{p^{2}-m_{h}^{2}+i\epsilon}$ (34)
as normally used in standard computations. The mass $m_{h}$ is given in eq.(III.2) and can be obtained by solving the corresponding set of gap equations. From these results, we can see that the “phion” can decay into a number of Higgs particles.
At this stage, we can write the gap equations explicitly in the form, keeping Euclidean metric,
$m_{\phi}^{2}$ $=$ $3\lambda_{\phi}\int\frac{d^{4}p}{(2\pi)^{4}}\sum_{n}\frac{B_{n}}{p^{2}+m_{n}^{2}}-\beta v^{2}-\beta\int\frac{d^{4}p}{(2\pi)^{4}}\frac{1}{p^{2}+m_{h}^{2}}$
$m_{h}^{2}$ $=$ $3\lambda_{h}v^{2}+3\lambda_{h}\int\frac{d^{4}p}{(2\pi)^{4}}\frac{1}{p^{2}+m_{h}^{2}}-\beta\int\frac{d^{4}p}{(2\pi)^{4}}\sum_{n}\frac{B_{n}}{p^{2}+m_{n}^{2}},$ (35)
where
$B_{n}=M_{\phi}{\hat{Z}}(m_{\phi},\lambda_{\phi})\frac{2\pi^{3}}{K^{3}(\kappa)}(-1)^{n}\frac{e^{-(n+\frac{1}{2})\pi\frac{K^{\prime}(\kappa)}{K(\kappa)}}}{1-e^{-(2n+1)\frac{K^{\prime}(\kappa)}{K(\kappa)}\pi}}(2n+1)^{2},$ (36)
and
$m_{n}(m_{\phi})=(2n+1)\frac{\pi}{2K(\kappa)}\sqrt{m_{\phi}^{2}+\frac{\lambda_{\phi}\mu^{4}}{m_{\phi}^{2}+\sqrt{m_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}},$ (37)
with
$\kappa^{2}=\frac{m_{\phi}^{2}-\sqrt{m_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}{m_{\phi}^{2}+\sqrt{m_{\phi}^{4}+2\lambda_{\phi}\mu^{4}}}.$ (38)
These equations can be solved very easily if we assume $\beta$ so small to give negligible contributions to these gap equations. For our convenience, we assume $\beta v^{2}$ finite and retain it. Therefore,
$\int\frac{d^{4}p}{(2\pi)^{4}}\sum_{n}\frac{B_{n}}{p^{2}+m_{n}^{2}}$ $=$ $-\frac{1}{16\pi^{2}}\sum_{n=0}^{\infty}\frac{\pi^{3}}{4K^{3}(i)}(2n+1)^{2}\frac{e^{-\left(n+\frac{1}{2}\right)\pi}}{1+e^{-(2n+1)\pi}}m_{n}^{2}(0),$
$\int\frac{d^{4}p}{(2\pi)^{4}}\frac{1}{p^{2}+m_{h}^{2}}$ $=$ $-\frac{1}{16\pi^{2}}m_{h}^{2}.$ (39)
Here, we have assumed the first iterate with the mass shift for the $\phi$ field is zero, and besides, the cut-off terms have been re-absorbed into the coupling constants $\lambda_{\phi}$ and $\lambda_{h}$. Therefore, we can finally approximate
$m^{2}_{\phi}$ $=$ $-\frac{3\lambda_{\phi}}{256}\frac{\pi^{3}}{K^{5}(i)}\sum_{n=0}^{\infty}(2n+1)^{4}\frac{e^{-\left(n+\frac{1}{2}\right)\pi}}{1+e^{-(2n+1)\pi}}\sqrt{\frac{\lambda_{\phi}}{2}}\mu^{2}-\beta v^{2},$
$m^{2}_{h}$ $=$ $3\lambda_{h}v^{2}-\frac{3\lambda_{h}}{16\pi^{2}}m_{h}^{2}.$ (40)
We are able to consistently solve this system of equations. E.g., for the Higgs mass one gets
$m_{h}^{2}=\frac{3\lambda_{h}v^{2}}{1+\frac{3\lambda_{h}}{16\pi^{2}}}.$ (41)
This is consistent with expectations. Taking into consideration the coupling $\beta$, we get the following set of equations
$m^{2}_{\phi}$ $=$ $-\frac{3\lambda_{\phi}}{16\pi^{2}}\sum_{n=0}^{\infty}\frac{\pi^{3}}{4K^{3}(i)}(2n+1)^{2}\frac{e^{-\left(n+\frac{1}{2}\right)\pi}}{1+e^{-(2n+1)\pi}}m_{n}^{2}(0)-\beta v^{2}+\frac{\beta}{16\pi^{2}}m_{h}^{2}$
$m^{2}_{h}$ $=$ $3\lambda_{h}v^{2}-\frac{3\lambda_{h}}{16\pi^{2}}m_{h}^{2}+\frac{\beta}{16\pi^{2}}\sum_{n=0}^{\infty}\frac{\pi^{3}}{4K^{3}(i)}(2n+1)^{2}\frac{e^{-\left(n+\frac{1}{2}\right)\pi}}{1+e^{-(2n+1)\pi}}m_{n}^{2}(0).$ (42)
These become
$m^{2}_{\phi}$ $=$ $-\frac{3\lambda_{\phi}}{16\pi^{2}}\sum_{n=0}^{\infty}\frac{\pi^{3}}{4K^{3}(i)}(2n+1)^{4}\frac{e^{-\left(n+\frac{1}{2}\right)\pi}}{1+e^{-(2n+1)\pi}}m_{0}^{2}-\beta v^{2}+\frac{\beta}{16\pi^{2}}m_{h}^{2}$
$m^{2}_{h}$ $=$ $3\lambda_{h}v^{2}-\frac{3\lambda_{h}}{16\pi^{2}}m_{h}^{2}+\frac{\beta}{16\pi^{2}}\sum_{n=0}^{\infty}\frac{\pi^{3}}{4K^{3}(i)}(2n+1)^{4}\frac{e^{-\left(n+\frac{1}{2}\right)\pi}}{1+e^{-(2n+1)\pi}}m_{0}^{2}.$ (43)
Here, $m_{0}$ is the ground state from eq.(33) for the $\phi$ field. Let us introduce a constant,
$\xi=\sum_{n=0}^{\infty}\frac{\pi^{3}}{4K^{3}(i)}(2n+1)^{4}\frac{e^{-\left(n+\frac{1}{2}\right)\pi}}{1+e^{-(2n+1)\pi}}\approx 21.2231\ldots,$ (44)
and express the set of equations as
$m^{2}_{\phi}$ $=$ $-\frac{3\lambda_{\phi}}{16\pi^{2}}\xi m_{0}^{2}-\beta v^{2}+\frac{\beta}{16\pi^{2}}m_{h}^{2},$
$m^{2}_{h}$ $=$ $3\lambda_{h}v^{2}-\frac{3\lambda_{h}}{16\pi^{2}}m_{h}^{2}+\frac{\beta}{16\pi^{2}}\xi m_{0}^{2}.$ (45)
The leading order solutions are given above can be written down as
${\bar{m}}^{2}_{\phi}$ $=$ $-\frac{3\lambda_{\phi}}{16\pi^{2}}\xi m_{0}^{2}-\beta v^{2},$
${\bar{m}}_{h}^{2}$ $=$ $\frac{3\lambda_{h}v^{2}}{1+\frac{3\lambda_{h}}{16\pi^{2}}}.$ (46)
This yields by iteration
$m_{\phi}^{2}$$\approx$ ${\bar{m}}^{2}_{\phi}+\frac{\beta}{16\pi^{2}}{\bar{m}}_{h}^{2},$
$M_{h}^{2}$$\approx$ ${\bar{m}}_{h}^{2}-\frac{\beta}{3\lambda_{\phi}}{\bar{m}}^{2}_{\phi},$ (47)
where we neglected $O(\beta^{2})$ terms. This implies $\beta/\lambda_{\phi}\ll 1$ to keep the mass shift for the $h$ field small. With this hypothesis, we can write the phion mass spectrum as
[…]
The $\mu$ parameter is critical for the physical consistency of the model. In Fig.1 we show how a set of parameters exists that yields a meaningful theoretical result with respect to experimental data.
Figure 1: We show the evolution of the phion mass $M_{\phi}=m_{0}$ with respect to $\mu$ and the corresponding evolution of the Higgs boson mass $M_{h}$ (blue dots) with the value of the experimental Higgs boson mass and its error bar. We set $\beta=10^{-4}$, $\lambda_{\phi}=10^{-2}$, $\lambda_{h}=0.086$ and $v=0.246\ \textit{TeV}$. The model appears to be consistent with a wide range of $\mu$ values.
We also show Fig.2, which is similar to Fig.1 but as a function of $\lambda_{\phi}$. We can achieve consistency.
Figure 2: Plot same as Fig.1 but as a function of $\lambda_{\phi}$ rather than $\mu$. It is seen that the consistency of the model is granted provided $\lambda_{\phi}$ is smaller enough. We assume $\mu=1.5\ TeV$.
We can get a general condition for the Higgs vacuum expectation value as it depends on the other parameters of the model. From eq.(III.1), we get, setting $H_{1}^{h}(x)=v$,
$m_{h}^{2}>\frac{\beta}{3\lambda_{h}}\xi m_{0}^{2}.$ (49)
We obtain the plot in Fig.3 for the inequality (49). The red curve is well below the Higgs boson mass for a large set of values of $\lambda_{\phi}$ as required and so, there exists a meaningful range of parameters for which the scenario is fully consistent.
Figure 3: In this plot, the limit of mass (blue line) is seen to be below the red line for a large set of $\lambda_{\phi}<1$, in agreement with our discussion in the text. This grants the full consistency of the model. We set $\mu=1.5\ TeV$.
IV Conclusions and Discussions
We investigated conformally extended Standard Model with a hidden scalar $\phi$ and showed that due to dynamics in the hidden sector with quartic potential (see eq.(5)), $\phi$ develops a vacuum expectation value (vev) in the form of a mass gap which triggers the electroweak symmetry breaking (EWSB). We summarise our main findings below:
  • We provide a novel pathway for dynamical generation of scales, particularly in the context of EW scale generation via dimensional transmutation from a hidden scalar sector starting from a scale-invariant theory at the classical level.
  • For this purpose we solved the Dyson-Schwinger Equations using the exact solution known via a novel technique developed, by Bender, Milton and Savage [8], working in the form of partial differential equations (see (III.1),(III.1) and (III.2)).
  • We derived analytically the Higgs boson mass which is dimensionally transmuted from the hidden sector shown in eq.(III.2).
  • This yields a consistent solution for the Higgs boson mass, in complete agreement with the experimental data, for a large set of the parameters of the theory for the given ordering. This is very well exemplified in the plots given in Fig.1, Fig.2 and Fig.3.
With null signatures of any SM extension at the LHC and in other searches, the framework of naturalness deserves to be re-examined. Among several ideas of explaining the EW scale dynamically generated, we discussed a scenario where conformal symmetry plays an essential role and the EW scale is a consequence of quantum effects just like QCD scale generation in the SM. We have shown that it is possible to generate the EW scale by including a new scalar which talks to the SM Higgs via a simple Higgs-portal coupling. The extension is rather minimal. The mass of this new scalar boson is constrained from the successful generation of the SM Higgs boson mass and the BSM microphysics parameters gets fixed within the uncertainties of the Higgs mass measurements.
There could be a way to search for the Higgs-portal scale in laboratories. Also, our scenario may have an impact on Higgs-portal dark matter model and some profound implications to EW phase transition in the early universe. We leave such investigations to future work.
V Acknowledgements
The work of NO is supported in part by the United States Department of Energy (DC-SC 0012447 and DC-SC 0023713).

### v2

@@
 However, from the observational point-of-view this mechanism fails within the Standard Model to generate the correct Higgs mass (the Electroweak (EW) Scale) as it predicts $m_{Z,W}>m_{H}$, where $m_{Z,W}$ are the SM Z and W gauge boson masses while $m_{H}$ is the SM Higgs boson mass [18, 22].
-Inspite of this faliure this has remained the direction of BSM model building and several extensions of the SM has been explored extensively in the literature [1, 17, 86, 18, 20, 19], where this radiative EW symmetry breaking (EWSB) mechanism is successful in terms of observations and often makes very concrete and testable predictions.
+In spite of this failure this has remained the direction of BSM model building and several extensions of the SM has been explored extensively in the literature [1, 18, 87, 19, 21, 20], where this radiative EW symmetry breaking (EWSB) mechanism is successful in terms of observations and often makes very concrete and testable predictions.
 When non-minimal coupling to gravity is introduced, such scenarios can provide naturally flat inflaton potentials [78, 75, 85, 87, 73, 72, 7, 90] and stable particle dark matter candidates [63, 76, 73, 74, 77, 6].
@@
 We assume $\lambda_{h}\ll 1$ and $\beta\ll\lambda_{\phi}$.
+Before to go through the full model, it would be helpful to understand the decoupled case with $\beta=0$.
+One is left with a quartic scalar theory.
+This case has been extensively studied by us in [40, 16] and two kinds of solutions are expected.
+The quartic model admits an exact solutions in terms of Jacobi elliptical functions for eq.(II) and all higher order correlation functions are expressed through them that is the hallmark of a Gaussian solution.
+On the other hand, as proven in [16], there is also the constant solution that arises from quantum fluctuations as in eq.(II) the term $G_{2}(0)$, when properly evaluated and regularized, yields a mass term.
+This term has the right sign to give rise to the Higgs mechanism as currently understood for the Standard Model.
+Thus, it is up to nature to decide what solution applies.
+For simplicity reasons, the Higgs mechanism appears the most economical one and seems the one observed in experiments.
+Indeed, the other solution implies an infinite tower of massive excitations that were not observed so far [^3] A comparison with the Coleman-Weinberg effective potential [18] could only be possible in the approximate limit of a small coupling while, in our case, we have an exact solution that holds at any value of the coupling itself..
 III.1 Dyson-Schwinger equations
@@
 where we neglected $O(\beta^{2})$ terms. This implies $\beta/\lambda_{\phi}\ll 1$ to keep the mass shift for the $h$ field small. With this hypothesis, we can write the phion mass spectrum as
-[…]
+$m_{n}=(2n+1)\frac{\pi}{2K(i)}\left(\sqrt[4]{\frac{\lambda_{\phi}}{2}}\mu+\frac{m_{\phi}^{2}}{2^{\frac{3}{2}}\sqrt[4]{\lambda_{\phi}}\mu}\right).$ (48)
 The $\mu$ parameter is critical for the physical consistency of the model.
@@
 We obtain the plot in Fig.3 for the inequality (49).
-The red curve is well below the Higgs boson mass for a large set of values of $\lambda_{\phi}$ as required and so, there exists a meaningful range of parameters for which the scenario is fully consistent.
+The red curve is well below the Higgs boson mass for a large set of values of $\lambda_{\phi}$ as required and so, there exists a meaningful range of parameters for which the scenario is fully consistent as we get the correctly observed EW Higgs mass observed at LHC, being $\lambda_{\phi}$ fixed (within the uncertainty of Higgs mass data), for $\mu\sim 1.5\ {\rm TeV}$ and the condition (49) granted.
 Figure 3: In this plot, the limit of mass (blue line) is seen to be below the red line for a large set of $\lambda_{\phi}<1$, in agreement with our discussion in the text.
@@
 The work of NO is supported in part by the United States Department of Energy (DC-SC 0012447 and DC-SC 0023713).
+Data sharing
+Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
+Conflict of interest
+The authors declare no conflict of interest.